What is the pH of a solution whose [H3O+] concentration is [1 x 10 -9]
1
Expert's answer
2020-12-07T07:43:43-0500

[H3O+] = 1×10-9


Solution:

We can convert between [H3O+] and pH using the following equation:

pH = - log[H3O+]

pH = - log(1×10-9) = 9

pH = 9


Answer: The pH of the solution is 9.